OpenStudy (anonymous):

The 48th term in the sequence 6, 13, 20, 27, ... is ??

OpenStudy (anonymous):

first term is \(6\) second term is \(6 + 7\) third term is \(6 + 7 + 7 = 6+2\times 7\) fourth term is \(6+3\times 7\)

OpenStudy (anonymous):

48th term is \(6+47\times 7\)

OpenStudy (texaschic101):

you could also use the formula

OpenStudy (anonymous):

yes, and the formula will give \(6+47\times 7\)

OpenStudy (anonymous):

what is the formula? and do i go ahead and find solve that? would my answer be 335?

OpenStudy (texaschic101):

an = a1 + (n-1) * d

OpenStudy (anonymous):

i think the formula is \(a_1+(n-1)d\) which in this case is \(6+47\times 7\)

OpenStudy (anonymous):

and yes, 335 is right

OpenStudy (texaschic101):

n = the number of the term you want to find a1 = the first number d = the common difference

OpenStudy (texaschic101):

it comes out the way satellite said

OpenStudy (anonymous):

i think i understand it now. Thank you guys so much!

OpenStudy (texaschic101):

YW :)